KINETIC SPECTROSCOPY OF NCl^{a}

\underline{ANATOLY V. KOMISSAROV}, MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Metastable NCl(a^{1}Δ) is a promising energy carrier for use in chemically driven iodine lasers. The present studies of NCl kinetics were conducted in support of efforts to develop an NCl(a)/I laser system. Photolysis of N_{3}Cl was used to generate NCl for spectroscopic and kinetic measurements. The kinetics of NCl in the ground and excited electronic states was studied by means of time-resolved direct absorption and emission measurements. Rate constants for removal of NCl(a) by O_{2}, H_{2}, Cl_{2}, I_{2}, ClN_{3} and self-annihilation (NCl(a)+NCl(a)) were determined. The reaction of NCl(a) with I_{2} and energy transfer from NCl(a) to I were examined using photolysis of N_{3}Cl/I_{2} and N_{3}Cl/CH_{2}I_{2} mixtures, respectively. Time-resolved fluorescence from I^{∗} was used to follow the reaction and transfer kinetics. Kinetic models were developed to simulate the I^{∗} time histories. The models, which were anchored by the NCl removal rate constants, were used to explore the kinetics of secondary photoproducts. The relevance of the results to NCl/I transfer lasers will be discussed.

aWork supported by Air Force Office of Scientific Research.